OpenStudy (r_josh10):

The area of the base of a prism is 21 cm2. The perimeter of the base is 20 cm. The height of the prism is 8 cm. What is the surface area of the prism? cm2

OpenStudy (r_josh10):

@iGreen @AnswerMyQuestions

OpenStudy (anonymous):

/

OpenStudy (igreen):

Oh..

OpenStudy (igreen):

A = 2b + ph Where 'b' is the area of the base, 'p' is the perimeter of the base, and 'h' is the height. Plug them in

OpenStudy (igreen):

A = 2(21) + (20)(8)

OpenStudy (igreen):

Simplify @r_josh10

OpenStudy (r_josh10):

ok um 202?

OpenStudy (igreen):

Oh, nevermind..that's correct.

OpenStudy (igreen):

2 * 21 = 42 20 * 8 = 160 160 + 42 = 202

OpenStudy (r_josh10):

2*21=42 + 20*8=160 160+48=202

OpenStudy (r_josh10):

thanks!

OpenStudy (igreen):

Np